OpenStudy (anonymous):

Can some help me with this? 11. Use the discriminant of the following quadratic equation to describe its solutions. x2 - 16x + 64 = 0 I got x = 8 A. There are no solutions. B. The solutions are real and distinct. C. The solutions are real and equal.* D. The solutions are complex.

OpenStudy (anonymous):

6. Use the discriminant of the following quadratic equation to describe its solutions. 3x^2 + 7x + 2 = 0 I got X = -2 X = -1/3 A. The solutions are real and distinct. B. The solutions are real and equal. C. The solutions are complex. D. There are no solutions.

OpenStudy (mertsj):

The problem says to use the discriminant. Do you know what the discriminant is?

OpenStudy (anonymous):

no, could you explain it please?

OpenStudy (mertsj):

The discriminant is the part of the quadratic equation that is under the radical: b^2-4ac

OpenStudy (anonymous):

Oh okay

OpenStudy (mertsj):

It determines the type of roots that you have.

OpenStudy (mertsj):

If it is 0 then you have the square root of 0 which is 0 so then you add 0 and subtract 0 which is the same thing so there is only 1 root and it is real.

OpenStudy (mertsj):

If the discriminant is positive then you have two roots which would be real.

OpenStudy (mertsj):

If the discriminant is negative then you have two roots which are complex numbers.

OpenStudy (anonymous):

for 11. I got C. and 6. I got B. am I incorrect?

OpenStudy (anonymous):

C for 6 sorry

OpenStudy (mertsj):

11. Real and equal 6. Real and distinct.

OpenStudy (anonymous):

Oh okay, thank you! do you mind if I ask you 3 more questions I wanna know if I'm right

OpenStudy (mertsj):

For number 11, the discriminant is 0 For number 6, the discriminant is 25

OpenStudy (mertsj):

0k

OpenStudy (anonymous):

(m^5)^3 (3sqrt(t))^2 k3/2

OpenStudy (anonymous):

I'm stuck on those 3 problems

OpenStudy (mertsj):

For the first one use this: (a^m)^n=a^(mn)

OpenStudy (mertsj):

For the second one: [3sqrt(t)]^2=3^2[sqrt(t)]^2

OpenStudy (mertsj):

For k3/2 that means 3k divided by 2 so you could write 1.5k